# Information Access of the Oppressed: A Problem-Posing Framework for Envisioning Emancipatory Information Access Platforms

### v1

@@ Abstract.
Online information access (IA) platforms, like other forms of media in our history, are targets of authoritarian capture. These concerns are particularly serious and urgent today in light of the rising levels of democratic erosion worldwide, the emerging capabilities of generative AI technologies such as AI persuasion, and the increasing concentration of economic and political power in the hands of Big Tech. This raises the question of what alternative IA infrastructure we must reimagine and build to mitigate the risks of authoritarian capture of our information ecosystems. We explore this question through the lens of Paulo Freire’s theories of emancipatory pedagogy. Freire’s theories provide a radically different lens for exploring IA’s sociotechnical concerns relative to the current dominating frames of fairness, accountability, confidentiality, transparency, and safety. We make explicit, with the intention to challenge, the dichotomy of how we relate to technology as either technologists—who envision and build technology—and its users. We posit that this mirrors the teacher-student relationship in Freire’s analysis. By extending Freire’s analysis to IA, we challenge the notion that it is the burden of the (altruistic) technologists to come up with interventions to mitigate the risks that emerging technologies pose to marginalized communities. Instead, we advocate that the first task for the technologists is to pose these as problems to the marginalized communities, to encourage them to make and unmake the technology as part of their material struggle against oppression. Their second task is to redesign our online technology stacks to structurally expose spaces for community members to co-opt and co-construct the technology in aid of their emancipatory struggles. We operationalize Freire’s theories to develop a problem-posing framework for envisioning emancipatory IA platforms of the future, and call on technologists to rid ourselves of our complicity in the global rise of authoritarianism and intentionally engage in building the infrastructure for resistance.
@@ 1. Motivation
On July 23, 2025, U.S. president Donald J. Trump issued an executive order titled “Preventing Woke AI in the Federal Government” (Trump, 2023) prohibiting federal procurement of AI models that incorporate concepts like “critical race theory, transgenderism, unconscious bias, intersectionality, and systemic racism; and discrimination on the basis of race or sex”. The executive order, while claiming to safeguard AI models from purported “ideological biases”, seeks to strong-arm AI companies to conform with the administration’s ideological agenda (Noble and Walsh, 2025). Likewise, China’s National Technical Committee 260 on Cybersecurity of SAC published a technical document [^1] https://www.tc260.org.cn/upload/2024-03-01/1709282398070082466.pdf mandating all Chinese generative AI models to not generate content that violates the country’s “core socialist values”. Consequently, the Chinese AI chatbot DeepSeek refuses to answer certain political questions, like on the topics of Tiananmen Square and Taiwanese independence (Lu, 2025). Such acts of ideological imposition on AI model outputs is also not limited only to state actors. In May 2025, xAI’s chatbot Grok promoted conspiracy theories about “white genocide” in South Africa (Kerr, 2025) after said conspiracy theories were mainstreamed by xAI’s owner Elon Musk (Imray, 2025). On another occasion the same year, Grok responded to users with antisemitic and pro-Nazi content (Taylor, 2025), which followed only a few months after Musk himself was embroiled in a controversy for making hand gestures at Trump’s presidential inauguration event that appeared to be Nazi salutes (Jacobsen, 2025).
@@
As AI becomes a widely used technology for information access (IA) (Chatterji et al., 2025), these instances of attempted capture can be seen to follow historical examples of regimes trying to capture and manipulate different forms of media (Dragomir and Horowitz, 2025), and fit into a growing trend of digital authoritarianism (Roberts and Oosterom, 2025)—defined as “the use of digital information technology by authoritarian regimes to surveil, repress, and manipulate domestic and foreign populations” (Polyakova and Meserole, 2019). Traditional media, social media, and IA platforms have historically been sites of conflict between oppression and emancipation. However, these concerns are particularly serious today in the context of: (i) Rising levels of democratic erosion (Gorokhovskaia and Grothe, 2024) and geopolitical conflicts (Taylor, 2023; United Nations Meetings Coverage and Press Releases, 2023), (ii) the increasing concentration of economic and political power in the hands of Big Tech (Oremus, 2017; Whittaker, 2021; Forum, 2024; Kak and West, 2023) which increasingly places itself in alignment with militarized and authoritarian actors (LaFrance, 2024; Lewis, 2025; Kanj, 2025; Fatafta, 2024; Akbari, 2025; Kang and Grauer, 2025; Kang, 2025), and (iii) emerging capabilities of generative AI technologies, such as AI persuasion (Burtell and Woodside, 2023; Carroll et al., 2023; Park et al., 2023; El-Sayed et al., 2024), and the potential for its weaponization within IA platforms for mass manipulation (Mitra et al., 2024) and generative propaganda (Daepp et al., 2025). The confluence of these factors calls for increased critical scrutiny from IA researchers on the evolving relationship between users and IA platforms (Mitra, 2025a), to quote: “Imagine every time you searched online or accessed information via your digital assistant, the information was presented to you exactly in the form most likely to alter your consumer preferences or political opinions. Or, consider AI-generated digital characters in ads and videos that appropriate marginalized identities to say or act in ways that real members of that community may be strongly opposed to—a new form of ‘Digital Blackface’ (Johnson and Joe, 2023).
@@
Even if we were to assume that IA companies take seriously their role as a pillar of civic society, the existence of such capabilities and infrastructure incentivizes authoritarian capture of platforms. There may be a temptation to search for algorithmic or policy interventions that can act as guardrails around AI technologies. But given the lack of transparency around how these models are trained by private corporations, identifying and enforcing such interventions become challenging. Also, many of the harms must be understood in the context of the IA platforms where these AI models are deployed and in context of their complex interactions with people searching for information. We must also not take it for granted that these AI technologies necessarily have a beneficial role in IA and must leave room for refusal of their deployment (Barocas et al., 2020). Finally, while AI capabilities may make IA platforms more attractive for capture, the risk does not go away simply by their exclusion from the platform. Therefore, we posit that the IA platforms are more meaningful sites for systemic mitigation. We ask: “What is the alternative IA infrastructure that we must reimagine and realize in practice as a form of mitigation for authoritarian capture of IA platforms?”.
@@
In the social media landscape, decentralization has emerged as a dominant strategy to safeguard against capture (Ciriello et al., 2025). Decentralization may also hold the key to protecting IA platforms—e.g., search engines. However, while the decentralization of technological infrastructure may be necessary, it is not sufficient to ensure that platforms are emancipatory (Hendrix and Flowers, 2022). Our pursuit for alternative information infrastructure is not value-neutral. In fact, our motivation stems from the recognition of IA as sites of political struggles between oppressors and oppressed. We reject any ambivalence or false neutrality with respect to oppressor-oppressed relations and situate our work in recent calls to explicitly align IA research with humanistic and emancipatory goals (Mitra, 2025a; Azzopardi et al., 2024; Trippas et al., 2025). To put it plainly, it is our emancipatory aspirations that prompt us to recognize the grave risks of state and corporate actors capturing our information platforms to manipulate mass opinions, and therefore it is necessary that we go beyond focusing on platform decentralization. This is in contrast to other frames of critique of information platform capture, like digital sovereignty (Djuric, 2025; FitzGerald, 2025) whose critique isn’t targeted at the practices of capture itself but focused on whose nationalistic interests are served by the platforms. Instead, in this work we develop a theoretical framework to envision IA platforms that foregrounds how oppressive power relations subverts knowledge production and consensus building, uplifts marginalized and oppressed voices, seeks epistemic justice (Fricker, 2007), and aims to realize humanistic and emancipatory futures for all globally.
@@
Practices of futuring (Fry, 2009) and envisioning (Reeves, 2012) shape technology development. Our collectively held sociotechnical imaginaries (Jasanoff and Kim, 2015) do not only imagine but also co-construct our digital futures (Mager and Katzenbach, 2021). In IA, these practices commonly manifest as scholarly publications (e.g., (Shah and Bender, 2024; Metzler et al., 2021)) and strategic workshops (e.g., (Trippas et al., 2025; Azzopardi et al., 2024; Hagen et al., 2026)). However, the dominant underlying values and epistemologies that shape these practices and their outcomes in the field often remain implicit (Vrijenhoek et al., 2023; Mitra, 2025b). Envisioning emancipatory IA poses two distinct challenges in this context. The first more obvious requirement is that the envisioning process must produce figurative blueprint for IA platforms that safeguard against authoritarian capture and contribute towards emancipatory societal outcomes. A second less obvious but nonetheless critical requirement is that the envisioning process must involve reflective participation of marginalized communities who should get to shape the values of the platform. This second requirement is motivated by insights drawn from critical scholarship on emancipatory pedagogy, in particular Paulo Freire’s canonical work “Pedagogy of the Oppressed” (Freire, 1970). In the context of emancipatory pedagogy, Freire emphasizes, that the oppressed are not simply targets but rather the executors of liberatory actions. He critiques what he refers to as the “banking concept of education” in which emancipatory knowledge is “gifted” by those committing the act of liberating to those who are purportedly being liberated. He writes:
@@
Attempting to liberate the oppressed without their reflective participation in the act of liberation is to treat them as objects which must be saved from a burning building; it is to lead them into the populist pitfall and transform them into masses which can be manipulated. … The conviction of the oppressed that they must fight for their liberation is not a gift bestowed by the revolutionary leadership, but the result of their own conscientização.
@@
When we embark to envision emancipatory IA platforms, we must vigorously reject the notion that it is the burden of a few individuals or institutions to determine the design that makes platforms emancipatory. Fundamentally, we must reject technosolutionism and the idea that there are some key elusive design tricks and system features that are waiting to be discovered that once incorporated renders IA platforms universally emancipatory and safe from capture in all diverse cultural and sociopolitical context. Rather the basis of emancipatory IA lies in holding space for reflective participation of community members in the process of envisioning, design, implementation, maintenance, and governance of platforms. This concept is central to the framework we propose in this work. The framework adopts Freire’s problem-posing approach by trying to articulate fundamental questions that emancipatory IA platforms must address. The framework intentionally stops there to make space for diverse reflection, dialogue, participation, collaboration, and negotiation by community members in the rest of the envisioning and actualization of emancipatory IA that we hope will follow our current work. Envisioned IA platforms must further create new opportunities for ongoing community participation in extending, maintaining, and governing platforms. Community participation is also necessary to critique and revise the problems posed by our framework. By foregrounding community participation this way, we want to move away from framings of IA where people relate to platforms as either technologists (those who envision, build, and maintain platforms) or users (those who use platforms); in favor of framings where we all see ourselves engaged in epistemic activities of knowledge production and consensus building through the making and unmaking of IA technologies.
The rest of this article is organized as follows: We next discuss some background concepts relevant to our framework in Section 2. As recent calls (Mitra, 2025a; Trippas et al., 2025) for aligning IA research with the goals of universal humanization and emancipation serves as the starting point of our work, we provide a brief introduction to emancipatory IA (§2.1) and then discuss other initiatives in the fields of IA and social media that are relevant to our work (§2.2). Next, we situate our work in Freire’s theories and analyses of emancipatory pedagogy in Section 3 to develop a critical perspective on emancipatory IA. In Section 4, we present our problem-posing framework for envisioning Emancipatory IA. In Section 5, we speculate about some of the inherent tensions involved in this work before finally concluding in Section 6.
2. Background
2.1. Emancipatory IA

### v2

@@ Abstract.
Online information access (IA) platforms are targets of authoritarian capture. We explore the question of how to safeguard our platforms while ensuring emancipatory outcomes through the lens of Paulo Freire’s theories of emancipatory pedagogy. Freire’s theories provide a radically different lens for exploring IA’s sociotechnical concerns relative to the current dominating frames of fairness, accountability, confidentiality, transparency, and safety. We make explicit, with the intention to challenge, the technologist-user dichotomy in IA platform development that mirrors the teacher-student relationship in Freire’s analysis. By extending Freire’s analysis to IA, we challenge the technologists-as-liberator frame where it is the burden of (altruistic) technologists to mitigate the risks of emerging technologies for marginalized communities. Instead, we advocate for Freirean Design (FD) whose goal is to structurally expose the platform for co-option and co-construction by community members in aid of their emancipatory struggles. Further, we employ Freire’s problem-posing approach within this framework to develop a method to envision future emancipatory IA platforms.
@@ 1. Introduction
On July 23, 2025, U.S. president Donald J. Trump issued an executive order titled “Preventing Woke AI in the Federal Government” (Trump, 2023) prohibiting federal procurement of AI models that incorporate concepts like “critical race theory, transgenderism, unconscious bias, intersectionality, and systemic racism; and discrimination on the basis of race or sex”. The executive order, while claiming to safeguard AI models from alleged “ideological biases”, seeks to strong-arm AI companies to conform with the Trump regime’s ideological agenda (Noble and Walsh, 2025). Likewise, China’s National Technical Committee 260 on Cybersecurity of SAC published a technical document [^1] https://www.tc260.org.cn/upload/2024-03-01/1709282398070082466.pdf mandating all Chinese generative AI models to not generate content that violates the country’s “core socialist values”. Consequently, the Chinese AI chatbot DeepSeek refuses to answer questions on certain political topics, like Tiananmen Square and Taiwanese independence (Lu, 2025). Such acts of ideological imposition on AI model outputs is not limited only to state actors. In May 2025, xAI’s chatbot Grok promoted conspiracy theories about “white genocide” in South Africa (Kerr, 2025) after said conspiracy theories were mainstreamed by xAI’s owner Elon Musk (Imray, 2025). On another occasion the same year, Grok responded to users with antisemitic and pro-Nazi content (Taylor, 2025), which followed only a few months after Musk himself was embroiled in a controversy for making hand gestures at Trump’s presidential inauguration event that appeared to be Nazi salutes (Jacobsen, 2025).
@@
As AI becomes a widely used technology for information access (IA) [^2] Information access refers to the focused interaction between people and information (Shah and Bender, 2024). Subfields of the study of IA include information retrieval (IR) and information filtering (IF). Search engines, recommender systems, and chatbots are common examples of IA systems. While social media may also facilitate access to information, their primary function is to support interactions between people and therefore are not typically considered as IA systems. (Chatterji et al., 2025), these instances of attempted capture emulate historical examples of regimes trying to manipulate different forms of media (Dragomir and Horowitz, 2025), and fit into a growing trend of digital authoritarianism (Roberts and Oosterom, 2025)—defined as “the use of digital information technology by authoritarian regimes to surveil, repress, and manipulate domestic and foreign populations” (Polyakova and Meserole, 2019). Traditional media, social media, and IA platforms have historically been sites of conflict between oppression and emancipation. However, these concerns are particularly serious today in light of: (i) Rising levels of democratic erosion (Gorokhovskaia and Grothe, 2024) and geopolitical conflicts (Taylor, 2023; United Nations Meetings Coverage and Press Releases, 2023), (ii) growing concentration of economic and political power in the hands of Big Tech (Oremus, 2017; Whittaker, 2021; Forum, 2024; Kak and West, 2023) which increasingly aligns itself with authoritarian actors (LaFrance, 2024; Lewis, 2025; Kanj, 2025; Fatafta, 2024; Akbari, 2025; Kang and Grauer, 2025; Kang, 2025), and (iii) emerging capabilities of generative AI, such as AI persuasion (Burtell and Woodside, 2023; Carroll et al., 2023; Park et al., 2023; El-Sayed et al., 2024), and the potential for its weaponization within IA platforms for mass manipulation (Mitra et al., 2024). As Mitra (2025a) warns: “Imagine every time you searched online or accessed information via your digital assistant, the information was presented to you exactly in the form most likely to alter your consumer preferences or political opinions. Or, consider AI-generated digital characters in ads and videos that appropriate marginalized identities to say or act in ways that real members of that community may be strongly opposed to—a new form of ‘Digital Blackface’ (Johnson and Joe, 2023).” The confluence of these factors calls for critical scrutiny from IA researchers on the evolving relationship between users and IA platforms.
@@
While AI capabilities may make IA platforms more attractive for capture, the risk of capture is salient to IA platforms more generally. Algorithmic or policy interventions around AI technologies alone are unlikely to produce sufficient guardrails; and safeguarding AI models from oppressive ideological imposition by their owners is inherently difficult given the lack of transparency around how these models are trained by private corporations. Instead, these risks are better analyzed and addressed in the context of the IA platforms where these models are deployed and in the context of their complex interactions with people and information. Centering the IA platform also crucially leaves room for refusal to deploy these models as a valid form of resistance (Barocas et al., 2020). Therefore, we ask: “What is the alternative IA infrastructure that we should reimagine and realize in practice as a form of mitigation for authoritarian capture of IA platforms?”.
@@
Our pursuit for alternative IA infrastructure is not value-neutral. Our motivation stems from the recognition of IA as sites of political struggles between oppression and emancipation. We reject any ambivalence or false neutrality with respect to oppressor-oppressed relations and situate our work in recent calls to explicitly align IA research with humanistic and emancipatory goals (Mitra, 2025a; Azzopardi et al., 2024; Trippas et al., 2025). It is our emancipatory aspirations that prompt us to recognize the grave risks of state and corporate actors capturing our information platforms to manipulate mass opinions. This is in contrast to other frames of critique of information platform capture, like digital sovereignty (Djuric, 2025; FitzGerald, 2025) that do not directly target the practices of capture, but is fueled by discontent about whose nationalistic interests are served by platforms. In contrast, we foreground the concerns of how oppressive power relations subverts knowledge production and consensus building, suppresses marginalized voices, and perpetuates epistemic injustices (Fricker, 2007). This distinction has practical implications for our work. For example, in the social media landscape, decentralization has emerged as a dominant strategy to safeguard against capture (Ciriello et al., 2025), which may also be key to protecting IA platforms. However, the decentralization of technological infrastructure alone is insufficient to ensure that platforms are emancipatory as they tend to reproduce dominant structures of oppression (Hendrix and Flowers, 2022). Instead, we need a broader reenvisioning of IA platforms.
@@
Practices of futuring (Fry, 2009) and envisioning (Reeves, 2012) shape technology development. Our collectively held sociotechnical imaginaries (Jasanoff and Kim, 2015) do not only imagine but also co-construct our digital futures (Mager and Katzenbach, 2021). In IA, these practices commonly manifest as scholarly publications (e.g., (Shah and Bender, 2024; Metzler et al., 2021)) and strategic workshops (e.g., (Trippas et al., 2025; Azzopardi et al., 2024; Hagen et al., 2026)). However, the dominant values and epistemologies that shape these visions often remain implicit (Vrijenhoek et al., 2023; Mitra, 2025b), and insulated from community purview. Here, envisioning emancipatory IA poses two distinct requirements. The first requirement is that the envisioning process should produce a figurative blueprint for IA platforms that safeguards against authoritarian capture and contribute towards emancipatory societal outcomes. A second less obvious but nonetheless critical requirement is that the envisioning process should involve reflective participation of marginalized communities, who should be integral to shaping the values of such platforms. We motivate this second requirement by drawing from critical scholarship on emancipatory pedagogy, in particular Paulo Freire’s canonical work “Pedagogy of the Oppressed” (Freire, 1970). In the context of emancipatory pedagogy, Freire emphasizes that the oppressed are not simply targets but rather the executors of liberatory actions. He critiques what he refers to as the “banking concept of education” in which emancipatory knowledge is “gifted” by those committing the act of liberating to those who are purportedly being liberated. He writes:
@@
  • Attempting to liberate the oppressed without their reflective participation in the act of liberation is to treat them as objects which must be saved from a burning building; it is to lead them into the populist pitfall and transform them into masses which can be manipulated… The conviction of the oppressed that they must fight for their liberation is not a gift bestowed by the revolutionary leadership, but the result of their own conscientização.
@@
When we embark to envision emancipatory IA platforms, we should vigorously reject the notion that it is the burden of a few individuals or institutions to determine the design that makes platforms emancipatory. We should also reject technosolutionism and the idea that there are some key elusive design tricks waiting to be discovered that once incorporated render IA platforms universally emancipatory and safe from capture in all diverse cultural and sociopolitical contexts. Rather, the basis of emancipatory IA lies in holding space for reflective participation of community members in the process of envisioning, design, implementation, maintenance, and governance of platforms. We operationalize this insight by conceptualizing a notion of Freirean design, and consequently exercise Freire’s problem-posing framework to articulate fundamental challenges that emancipatory IA platforms should address. We intentionally stop there to make space for diverse reflection, dialogue, participation, collaboration, and negotiation by community members in the rest of the envisioning and actualization of emancipatory IA that we hope will follow our work. We require proposals to address these challenges to further create new opportunities for active ongoing community participation. We foreground community participation in this way to move away from the modes of technology production where people relate to platforms as either technologists (those who envision, build, and maintain platforms) or users (those who use but cannot modify the platforms); in favor of futures where we all see ourselves engaged in epistemic activities of knowledge production and consensus building through the making and unmaking of IA technologies.
The rest of this article is organized as follows: We situate our contributions in the longstanding traditions of justice-oriented computing in Section 2. Next, we develop the idea of Freirean design drawing from Freire’s theories and analyses in Section 3. In Section 4, we pose a set of problems as the starting point for envisioning emancipatory IA. In Section 5, we speculate about the inherent tensions involved in this work before finally concluding in Section 6.
2. Justice-oriented computing
Several subfields of computing, like digital humanism, science and technology studies (STS), and human-computer interaction (HCI) have studied how computing intersects with concerns of justice and emancipation and how to reorient the field to support humanistic values and affect social justice. The Vienna manifesto on Digital humanism (Werthner, 2020) demands that “we must shape technologies in accordance with human values and needs, instead of allowing technologies to shape humans”; STS goes beyond studying the “origins, dynamics, and consequences of science and technology” to engage with questions of equity, politics, and social change (Barben et al., 2008); and HCI has cultivated a tradition of social justice research (Dombrowski et al., 2016; Chordia et al., 2024) and design activism (Thorpe, 2011; Fuad-Luke, 2013; Hauser et al., 2013) within its discipline. Salient to their concerns, is the question of what “better” futures we want to build for and the values that shape our process of envisioning and actualizing future technologies. Approaches such as value-sensitive design (Friedman, 1996) and speculative design (Auger, 2013) foreground these questions, but subsequently leave open the critical enquiry for what values (Albrechtslund, 2007; Manders-Huits, 2011; Reijers and Gordijn, 2019; Jacobs and Huldtgren, 2021) and “better” according to who (Bowen, 2010; Martins, 2014)?
2.1. Humanistic and emancipatory values in computing
To address the “value” question in computing, Asad (2019) proposed prefigurative design to explicitly reorient design work in service of “political or civic goals” and several other strands of research has explicitly grounded their work in liberatory epistemologies, including humanistic (Bardzell and Bardzell, 2016a, 2015), anti-oppressive and emancipatory (Smyth and Dimond, 2014; Bardzell and Bardzell, 2016a; Kane et al., 2021; Monroe-White, 2021; Saxena et al., 2023), feminist (Wajcman, 2004, 2010; Bardzell, 2010; Bardzell and Bardzell, 2011; Bardzell et al., 2011; Bardzell and Bardzell, 2016b; Bardzell, 2018; D’ignazio and Klein, 2020), queer (Light, 2011; Klipphahn-Karge et al., 2024; Guyan, 2022), postcolonial and decolonial (Irani et al., 2010; Philip et al., 2012; Dourish and Mainwaring, 2012; Sun, 2013; Ali, 2014, 2016; Akama et al., 2016; Irani and Silberman, 2016; Avle et al., 2017; Adams, 2021; Mohamed et al., 2020), anti-racist (Abebe et al., 2022; Ogbonnaya-Ogburu et al., 2020), indigenous (Winschiers-Theophilus and Bidwell, 2013; Shaw et al., 2014; Akama et al., 2016), anti-casteist (Kalyanakrishnan et al., 2018; Sambasivan et al., 2021; Vaghela et al., 2022a, b; Shubham, 2022; Kanjilal, 2023), anti-ableist (Williams et al., 2021; Sum et al., 2024), anti-fascist (McQuillan, 2022), abolitionist (Benjamin, 2019; Barabas, 2020; Earl, 2021; Jones and Melo, 2021; Williams and Haring, 2023), post-capitalistic (Feltwell et al., 2018; Browne and Green, 2022), and anarchist (Keyes et al., 2019; Linehan and Kirman, 2014; Asad et al., 2017) thought. Values-driven research have also been put to practice, e.g., the application of feminist HCI by Dimond et al. (Dimond, 2012; Dimond et al., 2013). These different strands of research are tied by their shared rejection of technological determinism (Héder, 2021) in favor of recognizing how technology and society mutually shape each other. It also calls for working towards alternative liberatory futures, aptly summarized by Keyes et al. (2019) as: “radically reorienting the field towards creating prefigurative counterpower—systems and spaces that exemplify the world we wish to see, as we go about building the revolution in increment”.